# Tensor Attention Training: Provably Efficient Learning of Higher-order Transformers

### v1

@@ 6.2 Main Results for Hardness
See the formal proof in Appendix G.2. The intuition is that if we can solve $\mathsf{ATAttLGC}$ in $O(t)$ time, then we can solve $\mathsf{ATAttC}$ in $O(t\cdot\log^{11}(n))$ time by interpolation and “integral”.
We see a similar sharp complexity transition as forward computation (Lemma 6.2), when we change the assumption from $B=O(\sqrt[3]{\log n})$ to $B=\Theta(\sqrt[3]{(1+\gamma)\log n})$ with $\gamma=\omega(1)$, where the tensor attention gradient computation will be unsolvable in truly sub-cubic time as well (assume $\mathsf{SETH}$).
7 Conclusion
In this work, we prove that the backward gradient of tensor attention training can be computed in almost linear $n^{1+o(1)}$ time, the same complexity as its forward computation, under a bounded entries assumption. We provide a closed-form solution for the gradient and propose a fast computation method utilizing polynomial approximation techniques and tensor algebraic tricks. Furthermore, we prove the necessity and tightness of our assumption through hardness analysis, showing that slightly weakening it renders the gradient problem unsolvable in truly subcubic time. Our theoretical results establish the feasibility of efficient higher-order transformer training and may facilitate practical applications of tensor attention architectures. Future work can explore how these findings can be implemented in real-world scenarios to enable the development of powerful multi-modality models.
@@ Acknowledgement
In Section A, we provide the potential limitations of this work. In Section B, we discuss the societal impacts of our work. In Section C, we provide general definitions and several basic facts. In Section D, we show how we calculate the gradient of the loss function defined by Definition 3.10. In Section E, we show the time complexity of our algorithm. In Section F, we show that our algorithm can be computed in polynomial time. In Section G, we show the hardness of our algorithm.
Appendix A Limitations
This work has not directly addressed the practical applications of our results, as we have not implemented our tensor attention algorithm. Future research could investigate how these findings could be applied in real-world scenarios.
Appendix B Societal Impacts
@@
We delve into and offer a deeper understanding of the attention mechanism, introducing a novel approach to integrate multi-modality into attention through the tensor attention algorithm. We also demonstrate that the computation of both forward and backward tensor attention can be achieved with almost linear time complexity.
Regarding the negative societal impact, since our work is completely theoretical in nature, we do not foresee any potential negative societal impacts which worth pointing out.
@@ Appendix C Tensor Operation Background
In Section C.1, we define the notation of computational time and the tensor operation. In Section C.2, we provide some helpful facts of tensor operation. In Section C.3, we provide some helpful facts of vectorization operation.
@@ C.1 General definitions and tensor operation
Definition C.3 ((,,)($\cdot$,$\cdot$,$\cdot$)( ⋅ , ⋅ , ⋅ )).
Let $\mathsf{X}\in\mathbb{R}^{d\times d\times d}$. Given matrices $A_{1}\in\mathbb{R}^{n\times d}$, $A_{2}\in\mathbb{R}^{n\times d}$ and $A_{3}\in\mathbb{R}^{n\times d}$.
Let operator $\mathsf{X}(A_{1},A_{2},A_{3})\in\mathbb{R}^{n\times n\times n}$ satisfying
$\mathsf{X}(A_{1},A_{2},A_{3})_{i,j,l}:=\sum_{a=1}^{d}\sum_{b=1}^{d}\sum_{c=1}^{d}\mathsf{X}_{a,b,c}(A_{1})_{i,a}(A_{2})_{j,b}(A_{3})_{l,c},~{}~{}~{}~{}\forall i\in[n],j\in[n],l\in[n]$
@@ C.2 Facts for tensor operation
Fact C.4.
We can show two results
@@
Fact C.8.
$\underbrace{(A_{1}\otimes A_{2})}_{{n^{2}\times d^{2}}}\underbrace{X^{\top}}_{{d^{2}\times d}}=\underbrace{A_{1}\oslash A_{2}}_{{n^{2}\times d}}$
Proof.
For any $i_{1},i_{2}\in[n],j\in[d]$, we have
$=$ $~{}(A_{1})_{i_{1},j}\cdot(A_{2})_{i_{2},j}$
$=$ $~{}(A_{1}\oslash A_{2})_{i_{1}+(i_{2}-1)n,j},$
Fact C.9 (Restatement of Fact 5.5).
@@ C.3 Facts for vectorization operation
Fact C.17 (Tensor-trick, Restatement of Fact 5.7).
@@
where the first step is due to the matrix being able to be written as a summation of vectors, the second step follows from Fact C.16, the third step follows from that matrix can be written as a summation of vectors, and the last step follows from the definition of vectorization operator $\operatorname{vec}(\cdot)$. ∎
Fact C.18.
@@
where the first step follows from Fact C.15, and the second step follows from Fact C.17. ∎
Fact C.19.
@@
where the first step follows from Fact C.18, the second step follows from the cyclic property of trace, the third step follows from Fact C.18, the fourth step follows from $A,B$ is symmetric, the fifth step is due to the definition of inner product, and the last step is due to Fact C.4.
@@ D.1 Definitions for useful functions
We define $A_{1},A_{2},A_{3}\in\mathbb{R}^{n\times d}$ to be three matrices in size $n\times d$. Suppose that $\mathsf{A}=A_{1}\otimes A_{2}\otimes A_{3}\in\mathbb{R}^{n^{3}\times d^{3}}$. Let $\mathsf{A}_{j_{0}}\in\mathbb{R}^{n^{2}\times d^{3}}$ represent an $n^{2}\times d^{3}$ sub-block from $\mathsf{A}$. There are $n$ such sub-blocks.
@@ D.3 Further information on gradient computation
$\frac{\mathrm{d}x}{\mathrm{d}x_{i}}=e_{i}$
@@ E.5 Closed form of gradient
From Fact 3.2, we know that
@@
where the first step comes from Definition 3.10, the second step is due to Eq. (E.5), the third step is because of Eq. (6), the fourth step is due to Eq. (7), the fifth step utilize the notation of $\operatorname{vec}(\cdot)$, and the last step follows from Fact C.17.
@@ F.1 Fast computation of $\mathsf{S}$
Thus, we can rewire $\mathsf{S}(x)=D^{-1}\exp(Q(K_{1}\oslash K_{2})^{\top}/d)\in\mathbb{R}^{n\times n^{2}}$ and we define $D=\operatorname{diag}(\exp(Q(K_{1}\oslash K_{2})^{\top}/d){\bf 1}_{n^{2}})$, where $Q=A_{1}X_{1},K_{1}=A_{2}X_{2},K_{2}=A_{3}X_{3}$.
@@ F.2 Fast computation of $\mathsf{V}$
$\|U_{1}(V_{1}\oslash W_{1})^{\top}h(y)-E-\mathsf{V}(x)\|_{\infty}=$ $~{}\|U_{1}(V_{1}\oslash W_{1})^{\top}\mathsf{L}(y)-E-\mathsf{S}(x)\mathsf{L}(y)+E\|_{\infty}$
@@ F.3 Fast computation of $\mathsf{W}$
According to Lemma F.2, we find a good approximation $U_{1}(V_{1}\oslash W_{1})^{\top}h(y)-E$ of $\mathsf{V}(x)$, where $k_{1}=n^{o(1)}$ and $U_{1},V_{1},W_{1}\in\mathbb{R}^{n\times k_{1}}$.
@@ F.4 Fast computation of $\mathsf{F}_{a}$: key step
where the first step is due to the definition of $\mathsf{F}_{a}(x)$, the second step is because of the definition of $U_{3},V_{3},W_{3}$, the third step is due to Fact C.9, the fourth step follows from the definition of $\widetilde{\mathsf{S}}(x)$ and $\widetilde{\mathsf{W}}(x)$, the fifth step is because of basic algebra, the sixth step comes from triangle inequality, and the last step is because bounded entries (we can write each number in $\mathsf{S}(x)$ and $\mathsf{W}(x)$ using $O(\log n)$ bits) and Lemma assumptions that $\|\widetilde{\mathsf{S}}(x)-\mathsf{S}(x)\|_{\infty}\leq\epsilon/\operatorname{poly}(n)$ and $\|\widetilde{\mathsf{W}}(x)-\mathsf{W}(x)\|_{\infty}\leq\epsilon/\operatorname{poly}(n)$
@@ F.5 Fast computation of $\mathsf{F}_{b}$: key step
It is noted that a good approximation of $\mathsf{S}(x)_{j_{0}}$ is $(U_{1}(V_{1}\oslash W_{1})^{\top})_{j_{0},*}^{\top}$ is . We denote the approximation of $\mathsf{S}(x)$ to be $\widetilde{\mathsf{S}}(x)=U_{1}(V_{1}\oslash W_{1})^{\top}$.
It is noted that a good approximation of $\mathsf{W}(x)_{j_{0}}$ is $(U_{2}(V_{2}\oslash W_{2})^{\top})_{j_{0},*}^{\top}$ is . Let denote the approximation of $\mathsf{W}(x)$ to be $\widetilde{\mathsf{W}}(x)=U_{2}(V_{2}\oslash W_{2})^{\top}$.
@@
For the side of computation time, we compute $V_{1}^{\top}V_{2}$ first and this takes $n^{1+o(1)}$ time.
@@
We can use $\widetilde{\mathsf{S}}(x)$ and $\widetilde{\mathsf{R}}(x)$ to approximate $\mathsf{F}_{b}(x)$. Let $\widetilde{\mathsf{F}}_{b}(x)=\underbrace{\operatorname{diag}(\widetilde{\mathsf{R}}(x))}_{n\times n}\underbrace{\widetilde{\mathsf{S}}(x)}_{n\times n^{2}}$. Because $\operatorname{diag}(\widetilde{\mathsf{R}}(x))$ is a diagonal matrix and $\widetilde{\mathsf{S}}(x)$ has low rank representation, then obviously we know how to construct $U_{4},V_{4},W_{4}$. Basically $U_{4}=\operatorname{diag}(\widetilde{\mathsf{R}}(x))U_{1}$ and $V_{4}=V_{1}$, $W_{4}=W_{1}$.
Now, we need to control the error, we have

### v2

@@ 6.2 Main Result for Hardness
See the formal proof in Appendix G.2. The intuition is that if we can solve $\mathsf{ATAttLGC}$ in $O(t)$ time, then we can solve $\mathsf{ATAttC}$ in $O(t\cdot\log^{11}(n))$ time by interpolation and “integral”. We see a similar sharp complexity transition as forward computation (Lemma 6.2): assuming $\mathsf{SETH}$, if we slightly weaken the assumption from $B=O(\sqrt[3]{\log n})$ to $B=\Theta(\sqrt[3]{(1+\gamma)\log n})$ with $\gamma=\omega(1)$, then the tensor attention gradient computation will be unsolvable in truly sub-cubic time as well.
7 Discussion and Conclusion
In this work, we proved that the backward gradient of tensor attention training can be computed in almost linear $n^{1+o(1)}$ time, the same complexity as its forward computation, under a bounded entries assumption. We provided a closed-form solution for the gradient and proposed a fast computation method utilizing polynomial approximation and tensor algebraic techniques. Furthermore, we proved the necessity and tightness of our assumption through hardness analysis, showing that slightly weakening it renders the tensor attention gradient problem unsolvable in truly subcubic time. Our theoretical results establish the feasibility of efficient higher-order transformer training and may facilitate practical applications of tensor attention architectures. Due to space limits, we provide our further discussion and extension in Appendix A. Future work can perform empirical evaluations of the method in practical large language models, and explore how these findings can be implemented in real-world scenarios to enable the development of powerful higher-order models.
@@ Acknowledgement
In Section A, we provide a further discussion and extension of this work. In Section B, we provide more related works. In Section C, we provide general definitions and several basic facts. In Section D, we show how we calculate the gradient of the loss function. In Section E, we show the time complexity of our algorithm. In Section F, we show that our algorithm can be computed in polynomial time. In Section G, we show the hardness of our algorithm.
Appendix A Further Discussion and Extension
Connection to real applications.
There are some empirical studies attempting to implement similar tensor attention (three order) in language modeling [81] and 3D medical image segmentation [116]. However, due to cubic time complexity, their models remain relatively small, e.g, 12M parameters in [81]. Although small scale, [81, 116] demonstrates the significant potential of tensor attention. Our work proves that an almost linear time algorithm for tensor attention mechanisms exists (Algorithm 1). This advancement could enable the scaling up of tensor attention and facilitate novel model designs in multi-modality, 3D imaging, and beyond. On the other hand, we abstract the most challenging part (the highest time complexity operation) in high-order attention into a clear mathematical problem and provide a solution. Our work introduces a new concept to the community, suggesting that cubic time complexity may not be the bottleneck in implementing three-order attention during training. Practical implementation poses additional significant challenges, considering numerous other techniques and operations, such as dropout, layer normalization, residual connections, position encoding, and many others. We hope our work inspires further algorithmic design.
Feasibility when the large value exists in the matrices.
If there exist many large entries in $Q,K_{1},K_{2},V_{1},V_{2}$, our hardness results (Theorem 6.3) indicate that no algorithm can accelerate the attention computation. However, several exciting works [92, 41] have shown that large entries are very sparse in the attention matrix. This suggests that our Algorithm 1 could inspire many potential practical implementations. One straightforward approach is to handle large entries separately, as in [41], and apply our algorithm to the remaining parts. There is undoubtedly a broad algorithm design space, and we hope our work provides valuable insights.
Extend our technique to compute the module-wise gradient.
Let $n$ be the input toke length, and $d$ be the hidden dimension. At the $i$-th layer of transformer model, let $G_{i}\in\mathbb{R}^{n\times d}$ denote the output of upstream gradient, $X_{i}\in\mathbb{R}^{n\times d}$ be defined in Definition 3.8, and $\mathsf{Attn}_{i}:=D^{-1}AV$ be the tensor attention model where $D,A,V$ are defined in Definition 3.5. Let $\mathsf{Loss}$ be some loss function. Then, by the chain rule, we have the module-wise gradient $\frac{\mathrm{d}\mathsf{Loss}}{\mathrm{d}X_{i}}=\mathrm{vec}(G_{i})\frac{\mathrm{d}\mathsf{Attn}_{i}}{\mathrm{d}X_{i}}$.
Extend our technique to the multi-head attention.
The gradient computation for each attention head in the same layer is independent of the others; each head only depends on its upstream gradient and its current module-wise gradient according to the chain rule. Therefore, our analysis can be directly applied to multi-head attention.
Generalize to scenarios involving multiple modalities
In our three-order attention, one attention module can handle three modalities simultaneously, i.e., $Q,K_{1},K_{2}$. For more modality, e.g., $m>3$ modality, there are two potential solutions in our minds. First, we could use $m$-order attention, i.e., $Q,K_{1},K_{2},\dots,K_{m-1}$. The inference and training time complexity for this approach are still unknown, and we leave it as our future work. Second, we could use multiple modules of three-order attention. Note that one layer of standard attention may introduce one more modality $K_{1}$ each time, while one layer of three-order attention may introduce two more modalities $K_{1},K_{2}$ each time. Thus, if we have $m+1$ modality and $Q$ is from one modality, say text, then the standard attention may need $m$ layers to merge all modalities together, whereas three-order attention may only need $\log(m)$ layers to merge them all together.
Societal impacts.
@@ Appendix A Further Discussion and Extension
We delve into and offer a deeper understanding of the attention mechanism, introducing a novel approach to integrate multi-modality into attention through the tensor attention algorithm. We also demonstrate that the computation of both forward and backward tensor attention can be achieved with almost linear time complexity. Regarding the negative societal impact, since our work is completely theoretical in nature, we do not foresee any potential negative societal impacts which worth pointing out.
Appendix B More Related Work
Large language models and transformer.
@@ Appendix C Tensor Operation Background
In Section C.1, we define the notation of computational time and the tensor operation. In Section C.2, we provide some helpful facts of tensor operation. In Section C.3, we provide some helpful facts of vectorization operation. In Section C.4, we provide some helpful facts about the tensor product.
@@ C.1 General definitions and tensor operation
We define the third mode tensor product, which is the core operator of tensor operations.
Definition C.2 (Third mode tensor product (,,)($\cdot$,$\cdot$,$\cdot$)( ⋅ , ⋅ , ⋅ )).
Let $\mathsf{X}\in\mathbb{R}^{d\times d\times d}$. Given matrices $A_{1}\in\mathbb{R}^{n\times d}$, $A_{2}\in\mathbb{R}^{n\times d}$ and $A_{3}\in\mathbb{R}^{n\times d}$. Let operator $\mathsf{X}(A_{1},A_{2},A_{3})\in\mathbb{R}^{n\times n\times n}$ satisfying
$\mathsf{X}(A_{1},A_{2},A_{3})_{i,j,l}:=\sum_{a=1}^{d}\sum_{b=1}^{d}\sum_{c=1}^{d}\mathsf{X}_{a,b,c}(A_{1})_{i,a}(A_{2})_{j,b}(A_{3})_{l,c},~{}~{}~{}~{}\forall i\in[n],j\in[n],l\in[n].$
@@ C.2 Facts for tensor operation
Fact C.4 (Transpose rule).
We show the results below
@@
Fact C.8 (Restatement of Fact 5.5).
@@ C.3 Facts for vectorization operation
Fact C.16 (Tensor-trick, Restatement of Fact 5.7).
@@
where the first step is due to the matrix being able to be written as a summation of vectors, the second step follows from Fact C.15, the third step follows from that matrix can be written as a summation of vectors, and the last step follows from the definition of vectorization operator $\operatorname{vec}(\cdot)$. ∎
Fact C.17.
@@
where the first step follows from Fact C.14, and the second step follows from Fact C.16. ∎
Fact C.18.
@@
where the first step follows from Fact C.17, the second step follows from the cyclic property of trace, the third step follows from Fact C.17, the fourth step follows from $A,B$ is symmetric, the fifth step is due to the definition of inner product, and the last step is due to Fact C.4. ∎
C.4 Facts for tensor product
Fact C.19.
$\underbrace{(A_{1}\otimes A_{2})}_{{n^{2}\times d^{2}}}\underbrace{X^{\top}}_{{d^{2}\times d}}=\underbrace{A_{1}\oslash A_{2}}_{{n^{2}\times d}}.$
Proof.
For any $i_{1},i_{2}\in[n],j\in[d]$, we have
$=$ $~{}(A_{1})_{i_{1},j}\cdot(A_{2})_{i_{2},j}$
$=$ $~{}(A_{1}\oslash A_{2})_{i_{1}+(i_{2}-1)n,j},$
@@ C.4 Facts for tensor product
Now, we introduce a key claim that can reduce the tensor product to matrix multiplication and Kronecker product to make calculation easy.
Claim C.21.
Let $\mathsf{I}_{d}\in\mathbb{R}^{d\times d\times d}$ and $A_{1},A_{2},A_{3}\in\mathbb{R}^{n\times d}$. We have $\mathrm{mat}(\mathsf{I}_{d}(A_{1},A_{2},A_{3}))=A_{1}\mathrm{mat}(\mathsf{I}_{d})(A_{2}\otimes A_{3})^{\top}=A_{1}(A_{2}\oslash A_{3})^{\top}\in\mathbb{R}^{n\times n^{2}}$
Proof.
The proof follows from Claim C.20 and Fact C.19. ∎
@@ D.1 Definitions for useful functions
We define $A_{1},A_{2},A_{3}\in\mathbb{R}^{n\times d}$ to be three matrices in size $n\times d$. Suppose that $\mathsf{A}=A_{1}\otimes A_{2}\otimes A_{3}\in\mathbb{R}^{n^{3}\times d^{3}}$. Let $\mathsf{A}_{j_{0}}\in\mathbb{R}^{n^{2}\times d^{3}}$ represent an $n^{2}\times d^{3}$ sub-block from $\mathsf{A}$. There are $n$ such sub-blocks, i.e. the $(i+(j_{0}-1)\cdot n^{2})$-th row, $j$-th column of $\mathsf{A}$ is the $i$-th row, $j$-th column of $\mathsf{A}_{j_{0}}$, for $i\in[n^{2}],j\in[d^{3}],j_{0}\in[n]$.
@@ D.3 Further information on gradient computation
$\frac{\mathrm{d}x}{\mathrm{d}x_{i}}=$ $~{}\frac{\mathrm{d}[x_{1},x_{2},\dots,x_{d^{3}}]^{\top}}{\mathrm{d}x_{i}}$
$=$ $~{}e_{i}$
where the first step follows from $x$ is a vector, and the second step follows from all coordinates are independent to each other.
@@
where the second step follows from Part 1.
@@ E.5 Closed form of gradient
We know that for all $a,b\in\mathbb{R}^{n}$, we have $\mathrm{diag}(a)\cdot b=\mathrm{diag}(b)\cdot a=a\circ b=b\circ a$. Then, we have
@@
where the first step comes from Definition 3.8, the second step is due to Eq. (E.5), the third step is because of Eq. (6), the fourth step is due to Eq. (7), the fifth step utilize the notation of $\operatorname{vec}(\cdot)$, and the last step follows from Fact C.16.
@@ F.1 Fast computation of $\mathsf{S}$
Thus, we can rewrite $\mathsf{S}(x)=D^{-1}\exp(Q(K_{1}\oslash K_{2})^{\top}/d)\in\mathbb{R}^{n\times n^{2}}$ and we define $D=\operatorname{diag}(\exp(Q(K_{1}\oslash K_{2})^{\top}/d){\bf 1}_{n^{2}})$, where $Q=A_{1}X_{1},K_{1}=A_{2}X_{2},K_{2}=A_{3}X_{3}$.
@@ F.2 Fast computation of $\mathsf{V}$
$\|U_{1}(V_{1}\oslash W_{1})^{\top}\mathsf{L}(y)-E-\mathsf{V}(x)\|_{\infty}=$ $~{}\|U_{1}(V_{1}\oslash W_{1})^{\top}\mathsf{L}(y)-E-\mathsf{S}(x)\mathsf{L}(y)+E\|_{\infty}$
@@ F.3 Fast computation of $\mathsf{W}$
According to Lemma F.2, we find a good approximation $U_{1}(V_{1}\oslash W_{1})^{\top}\mathsf{L}(y)-E$ of $\mathsf{V}(x)$, where $k_{1}=n^{o(1)}$ and $U_{1},V_{1},W_{1}\in\mathbb{R}^{n\times k_{1}}$.
@@ F.4 Fast computation of $\mathsf{F}_{a}$: key step
where the first step is due to the definition of $\mathsf{F}_{a}(x)$, the second step is because of the definition of $U_{3},V_{3},W_{3}$, the third step is due to Fact C.8, the fourth step follows from the definition of $\widetilde{\mathsf{S}}(x)$ and $\widetilde{\mathsf{W}}(x)$, the fifth step is because of basic algebra, the sixth step comes from triangle inequality, and the last step is because bounded entries (we can write each number in $\mathsf{S}(x)$ and $\mathsf{W}(x)$ using $O(\log n)$ bits) and Lemma assumptions that $\|\widetilde{\mathsf{S}}(x)-\mathsf{S}(x)\|_{\infty}\leq\epsilon/\operatorname{poly}(n)$ and $\|\widetilde{\mathsf{W}}(x)-\mathsf{W}(x)\|_{\infty}\leq\epsilon/\operatorname{poly}(n)$
@@ F.5 Fast computation of $\mathsf{F}_{b}$: key step
It is noted that a good approximation of $\mathsf{S}(x)_{j_{0}}$ is $(U_{1}(V_{1}\oslash W_{1})^{\top})_{j_{0},*}^{\top}$. We denote the approximation of $\mathsf{S}(x)$ to be $\widetilde{\mathsf{S}}(x)=U_{1}(V_{1}\oslash W_{1})^{\top}$.
It is noted that a good approximation of $\mathsf{W}(x)_{j_{0}}$ is $(U_{2}(V_{2}\oslash W_{2})^{\top})_{j_{0},*}^{\top}$. Let denote the approximation of $\mathsf{W}(x)$ to be $\widetilde{\mathsf{W}}(x)=U_{2}(V_{2}\oslash W_{2})^{\top}$.
@@
For the side of computation time, we compute $V_{1}^{\top}V_{2}$ first and this takes $n^{1+o(1)}$ time. Then, we compute $W_{1}^{\top}W_{2}$ and this also takes $n^{1+o(1)}$ time.
@@
We can use $\widetilde{\mathsf{S}}(x)$ and $\widetilde{\mathsf{R}}(x)$ to approximate $\mathsf{F}_{b}(x)$. Let $\widetilde{\mathsf{F}}_{b}(x)=\underbrace{\operatorname{diag}(\widetilde{\mathsf{R}}(x))}_{n\times n}\underbrace{\widetilde{\mathsf{S}}(x)}_{n\times n^{2}}$. Because $\operatorname{diag}(\widetilde{\mathsf{R}}(x))$ is a diagonal matrix and $\widetilde{\mathsf{S}}(x)$ has low-rank representation, then obviously we know how to construct $U_{4},V_{4},W_{4}$. Basically $U_{4}=\operatorname{diag}(\widetilde{\mathsf{R}}(x))U_{1}$ and $V_{4}=V_{1}$, $W_{4}=W_{1}$.
Now, we need to control the error, and we have